1. (10 pts.) Find the equations of the two tangent lines to the parametric curve r = y = 20 - 81 at the point (4,0).
Answer #1

Curves (4,0) x=42 ry= 245_st3 let us find the logo of the tangent, line de for that differentiate kly wanit ť die at TH Y= o3 x ²6 x dr. and point (410) ра) 20 Море . 14, - 1 point-slope forur of a live Yay, = m(x-x) 7-р = t (x-1) 6. с У-у, з мя- )